OpenStudy (anonymous):

Mr. Jones has taken a survey of college students and found that 1 out of 7 students are liberal arts majors. If a college has 12,000 students, what is the best estimate of the number of students who are liberal arts majors

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

1 out of 7 students are liberal arts majors so the ratio is 1/7 if a college has 12000 students -> 1/7 = x/12000 x=12000/7=1714.28 , x approximates 1714 x is the number of students who are liberal arts major